# Subspaces of $L^{2}(\mathbb{R}^{n})$ Invariant Under Shifts by a Crystal Group

### v1

@@ Abstract
For a crystal group $\Gamma$ in dimension $n$, a closed subspace $\mathcal{V}$ of $L^{2}(\mathbb{R}^{n})$ is called $\Gamma$–shift invariant if, for every $f\in\mathcal{V}$, the shifts of $f$ by every element of $\Gamma$ also belong to $\mathcal{V}$. The main purpose of this paper is to provide a characterization of the $\Gamma$–shift invariant closed subspaces of $L^{2}(\mathbb{R}^{n})$.
@@ 1 Introduction
Let $\Gamma$ be a crystal group in dimension $n$ with point group $\Pi$ and associated lattice $\mathcal{L}$. Detailed definitions are given in Section 3. Elements of $\Gamma$ are written $[x,M]$, where $M$ is an orthogonal matrix that belongs to $\Pi$ and $x$ is a vector in $\mathbb{R}^{n}$. For a function $f$ on $\mathbb{R}^{n}$, the shift of $f$ by $[x,M]$ is $\pi[x,M]f$ given by
@@
for $z\in\mathbb{R}^{n}$. When these shifts are applied to functions in the Hilbert space of all square-integrable functions, $L^{2}(\mathbb{R}^{n})$, they constitute a unitary representation of $\Gamma$, which we call the natural representation. A closed subspace $\mathcal{V}$ of $L^{2}(\mathbb{R}^{n})$ is called $\pi$-invariant, or $\Gamma$–shift invariant, if $\pi[x,M]f\in\mathcal{V}$, for every $f\in\mathcal{V}$ and every $x,M]\in\Gamma$. Our purpose is to provide a characterization of the $\Gamma$–shift invariant closed subspaces of $L^{2}(\mathbb{R}^{n})$.
@@
Another reason to be concerned about the nature of $\Gamma$–shift invariant closed subspaces is the emergence of topological quantum chemistry (see [8] and [7], for example) where band representations play a significant role. A band representation of a crystal group $\Gamma$ is the restriction of the natural representation $\pi$ to a closed $\Gamma$–shift invariant subspace generated by a single function in $L^{2}(\mathbb{R}^{n})$. A readable review of band representations can be found in [1].
@@
Our characterization of $\Gamma$–shift invariant closed subspaces of $L^{2}(\mathbb{R}^{n})$ is stated in Theorem 6.3 with a variation given in Corollary 6.7. Imprecisely stated, there is a one-to-one correspondence between $\Gamma$–shift invariant closed subspaces of $L^{2}(\mathbb{R}^{n})$ and certain maps (the range functions) from an open subset of $\mathbb{R}^{n}$ and the closed subspaces of $\ell^{2}(\mathcal{L}^{*})$, where $\mathcal{L}^{*}$ is the dual lattice of $\mathcal{L}$. If $\Gamma=\mathbb{Z}^{n}$, then the characterization in Theorem 6.3 reduces to the characterization of shift-invariant subspaces given in Proposition 1.5 in [5]. Our approach draws on the presentation of range functions by Bownik and Ross in [6], to the extent that $\Gamma$–shift invariance implies invariance under shifts by the lattice $\mathcal{L}$. Perhaps the feature of our characterization that is most interesting arises when $\Gamma$ is non-symmorphic; that is $\Gamma$ is not a semi-direct product of $\mathcal{L}$ with the point group $\Pi$. We view $\pi$ as a unitary representation of $\Gamma$ and our strategy is to use unitary maps to transform $\pi$ into an equivalent representation that can be analyzed using range function methods.
@@
We gather together preliminary facts and known results in Section 2 while Section 3 is devoted to organizing the properties of crystal groups that are needed. The unitary maps that will be used to transform $\pi$ are introduced in Section 4 and applied in the next section. In the final section, we prove our main theorem, Theorem 6.3. We also present an example of a non-symmorphic wallpaper group $\Gamma$ and a $\Gamma$–shift invariant closed subspace of $L^{2}(\mathbb{R}^{2})$ where the impact of the non-symmorphic nature of this $\Gamma$ is illustrated.
@@ 3 Crystal groups
Fix a cross-section $\gamma:\Pi\to\Gamma$ of the $T$-cosets in $\Gamma$; so $q\big(\gamma(M)\big)=M$, for all $M\in\Pi$. With $\gamma$ fixed, for each $M\in\Pi$, let $x_{M}\in\mathbb{R}^{n}$ be such that $\gamma(M)=[x_{M},M]$. Then $\Gamma=\{[\ell+x_{M},M]:\ell\in\mathcal{L},M\in\Pi\}$. For many crystal groups, $0,M]\in\Gamma$, for all $M\in\Pi$. Then, choose $x_{M}=0$, for each $M\in\Pi$. When this can be done, $\Gamma$ is isomorphic to the semidirect product $\mathcal{L}\rtimes\Pi$. Such crystal groups are called symmorphic.
@@
With the $x_{M}\in\mathbb{R}^{n}$ selected as above, $N^{-1}x_{K}+x_{N}-x_{KN}\in\mathcal{L}$, for all $K,N\in\Pi$.
@@
This auxilliary group might be isomorphic to $\Gamma$, but this is not always the case. As with $\Gamma$, $\mathcal{L}^{*}\rtimes\Pi$ acts on $\mathbb{R}^{n}$. For $(\nu,M)\in\mathcal{L}^{*}\rtimes\Pi$ and $\omega\in\mathbb{R}^{n}$, let $(\nu,M)\cdot\omega=M(\omega+\nu)$. This identifies $\mathcal{L}^{*}\rtimes\Pi$ with a discrete group of isometries of $\mathbb{R}^{n}$ such that, since $\mathcal{L}^{*}$ is a full-rank lattice, $\mathbb{R}^{n}/(\mathcal{L}^{*}\rtimes\Pi)$ is compact. That is, $\mathcal{L}^{*}\rtimes\Pi$ is also a crystal group. Let $\Gamma^{*}=\mathcal{L}^{*}\rtimes\Pi$.
@@ 4 Several Unitary Transformations
In order to illuminate aspects of the natural unitary representation of the crystal group $\Gamma$ in the next section, the Hilbert space $L^{2}(\mathbb{R}^{n})$ will be transformed into a number of different realizations by specific unitary maps. The first is the Fourier transform viewed as a unitary map of $L^{2}(\mathbb{R}^{n})$ onto itself.
@@
The function $w_{2}$ plays an important role in the characterization of $\Gamma$–shift invariant closed subspaces of $L^{2}(\mathbb{R}^{n})$ obtained in Section 6, so it is important to note that this function is independent of the choice of the cross-section $\gamma:\Pi\to\Gamma$. Indeed, if $\gamma^{\prime}$ is a different cross-section, then, for each $L\in\Pi$, $\gamma^{\prime}(L)=[x^{\prime}_{L},L]$, for some $x^{\prime}_{L}\in\mathbb{R}^{n}$. But $x_{L},L]^{-1}[x^{\prime}_{L},L]=[x^{\prime}_{L}-x_{L},{\rm id}]\in T$, so $k=x^{\prime}_{L}-x_{L}\in\mathcal{L}$. Thus, $e^{-2\pi i\nu\cdot x^{\prime}_{L}}=e^{-2\pi i\nu\cdot(x_{L}+k)}=e^{-2\pi i\nu\cdot x_{L}}$.
@@ 5 The natural representation
The action of $\Gamma$ on $\mathbb{R}^{n}$ generates a unitary representation. For $x,M]\in\Gamma$, $\pi[x,M]$ is the unitary operator given by, for $f\in L^{2}(\mathbb{R}^{n})$,
@@ 6 Invariant closed subspaces
The patch of brick wall illustrated in Figure 1 is meant to be a region of a pattern extending in all directions. The symmetry group of this pattern is often denoted $pg$, so set $\Gamma=pg$. If the origin is placed at the bottom of one of the vertical line segments, let $\bf u$} denote a horizontal vector pointing right whose length equals the length of one brick. Let $\bf v$} be an upward pointing vector whose length equals twice the width of a brick. Then, we can take the translation lattice of $\Gamma$ to be $\mathcal{L}=\{j{\bf u}+k{\bf v}:j,k\in\mathbb{Z}\}$. Let $\sigma\in{\rm O}_{2}$ denote the reflection about the horizontal axis. This is not a symmetry of the pattern, but $\frac{1}{2}{\bf u},\sigma]\in\Gamma$. Indeed, the point group of $\Gamma$ is $\Pi=\{{\rm id},\sigma\}$, isomorphic to $\mathbb{Z}_{2}$. If $T=\{[j{\bf u}+k{\bf v},{\rm id}]:j,k\in\mathbb{Z}\}$ is the translation subgroup of $\Gamma$, we can take the cross-section of the $T$-cosets to be given by $\gamma({\rm id})=[0,{\rm id}]$ and $\gamma(\sigma)=[\frac{1}{2}{\bf u},\sigma]$.
@@
Let’s construct a simple $\Pi$-invariant measurable range function for $\ell^{2}(\mathcal{L}^{*})$ based on $\Pi\Omega$. Let $\kappa=\begin{pmatrix}1\\ 0\end{pmatrix}\in\mathcal{L}^{*}$ and consider the one-dimensional subspace $V=\mathbb{C}(\delta_{0}+\delta_{\kappa})=\{\alpha(\delta_{0}+\delta_{\kappa}):\alpha\in\mathbb{C}\}$ of $\ell^{2}(\mathcal{L}^{*})$. Let $E$ be a measurable subset of $\Omega$ as pictured in Figure 3 where we have surpressed the vertical axis and changed the scale a little.
@@
which by Lemma 6.6 is equivalent to $G(L\omega)\in J^{\prime}(L\omega)$ for a.e. $\omega$, where $J^{\prime}$ is the $\Pi$-invariant extension of $J$. Thus $G\in\mathcal{M}_{J^{\prime}}$. It follows now from Theorem 6.3 that $\mathcal{V}$ is invariant under shifts by $\Gamma$. ∎

### v2

# arXiv:2601.11839v2[math.FA] 28 Feb 2026

@@ Abstract
For a crystal group $\Gamma$ in dimension $n$, a closed subspace $\mathcal{V}$ of $L^{2}(\mathbb{R}^{n})$ is called $\Gamma$–shift-invariant if, for every $f\in\mathcal{V}$, the shifts of $f$ by every element of $\Gamma$ also belong to $\mathcal{V}$. The main purpose of this paper is to provide a characterization of the $\Gamma$–shift-invariant closed subspaces of $L^{2}(\mathbb{R}^{n})$ that is valid for all crystal groups, non-symmorphic as well as symmorphic.
@@ 1 Introduction
Let $\Gamma$ be a crystal group in dimension $n$ with point group $\Pi$ and associated lattice $\mathcal{L}$. Detailed definitions are given in Section 3. Elements of $\Gamma$ are written $[x,M]$, where $M$ is an orthogonal matrix that belongs to $\Pi$ and $x$ is a vector in $\mathbb{R}^{n}$. The crystal group is called symmorphic if $0,M]\in\Gamma$, for all $M\in\Pi$; or, equivalently, if $\Gamma$ is isomorphic to the semi-direct product $\mathcal{L}\rtimes\Pi$. If this is not the case, then $\Gamma$ is non-symmorphic.
For a function $f$ on $\mathbb{R}^{n}$, the shift of $f$ by $[x,M]$ is $\pi[x,M]f$ given by
@@
for $z\in\mathbb{R}^{n}$. When these shifts are applied to functions in the Hilbert space of all square-integrable functions, $L^{2}(\mathbb{R}^{n})$, they constitute a unitary representation $\pi$ of $\Gamma$, which we call the natural representation. A closed subspace $\mathcal{V}$ of $L^{2}(\mathbb{R}^{n})$ is called $\pi$-invariant, or $\Gamma$–shift-invariant, if $\pi[x,M]f\in\mathcal{V}$, for every $f\in\mathcal{V}$ and every $x,M]\in\Gamma$. Our purpose is to provide a characterization of the $\Gamma$–shift-invariant closed subspaces of $L^{2}(\mathbb{R}^{n})$.
@@
Another reason to be concerned about the nature of $\Gamma$–shift-invariant closed subspaces is the emergence of topological quantum chemistry (see [9] and [8], for example) where band representations play a significant role. A band representation of a crystal group $\Gamma$ is the restriction of the natural representation $\pi$ to a closed $\Gamma$–shift-invariant subspace generated by a single function in $L^{2}(\mathbb{R}^{n})$. A readable review of band representations can be found in [1].
In 2020, Barbieri, et al, [3] considered $\Gamma$–shift-invariance of closed subspaces of $L^{2}(\mathcal{G})$, where $\mathcal{G}$ is a locally compact abelian group and $\Gamma$ is the semi-direct product, $\Lambda\rtimes G$, where $\Lambda$ is a uniform lattice in $\mathcal{G}$ and $G$ is a finite group of continuous automorphisms of $\mathcal{G}$ that fix $\Lambda$. Theorem 4.7 of [3] provides a characterization of $\Gamma$–shift-invariant closed subspaces of $L^{2}(\mathcal{G})$. When $\mathcal{G}=\mathbb{R}^{n}$, this generalizes the characterization of shift-invariant closed subspaces of $L^{2}(\mathbb{R}^{n})$ given in [6] to $\Gamma$–shift-invariance, where $\Gamma$ is a symmorphic crystal group.
Our main result, given in Theorem 6.3 and reformulated in Corollary 6.7, applies to all crystal groups. Thus, the contribution of this paper is to describe the situation for non-symmorphic crystal groups. The two crystal groups in one dimension are symmorphic, while 4 of the 17 two dimensional crystal groups are non-symmorphic. In three dimensions, 146 out of the 219 different (up to isomorphism) crystal groups are non-symmorphic. It gets more dramatic as the dimension increases. In dimensions 5 and 6, more than 85% and 97%, respectively, of the crystal groups are non-symmorphic (see [19] for exact numbers). A readable account of the process of enumeration of crystal groups is given in [18]. Also, Section 2.5 of [18] contains an explanation of the connection between a vector system for a given point group and the non-symmorphic crystal groups with that point group. We will discuss how a vector system can be defined for a particular $\Gamma$ in Section 3 and it turns out that the vector system plays a key role in our main result.
@@
Imprecisely stated, we show, in Theorem 6.3 and Corollary 6.7, that there is a one-to-one correspondence between $\Gamma$–shift-invariant closed subspaces of $L^{2}(\mathbb{R}^{n})$ and certain maps (twisted range functions, where the twist comes from the vector system) from an open subset of $\mathbb{R}^{n}$ and the closed subspaces of $\ell^{2}(\mathcal{L}^{*})$, where $\mathcal{L}^{*}$ is the dual lattice of $\mathcal{L}$. If $\Gamma=\mathbb{Z}^{n}$, then the characterization in Theorem 6.3 reduces to the characterization of shift-invariant subspaces given in Proposition 1.5 in [6]. When $\Gamma$ happens to be symmorphic, Theorem 6.3 is equivalent to the characterization for symmorphic crystal groups found in [3]. Our approach draws on the presentation of range functions by Bownik and Ross in [7], to the extent that $\Gamma$–shift-invariance implies invariance under shifts by the lattice $\mathcal{L}$. We view $\pi$ as a unitary representation of $\Gamma$ and our strategy is to use unitary maps to transform $\pi$ into an equivalent representation that can be analyzed using range function methods.
@@
We gather together preliminary facts and known results in Section 2 while Section 3 is devoted to organizing the properties of crystal groups that are needed. The unitary maps that will be used to transform $\pi$ are introduced in Section 4 and applied in the next section. In the final section, we prove our main theorem, Theorem 6.3. We also present an example of a non-symmorphic wallpaper group $\Gamma$ and a $\Gamma$–shift-invariant closed subspace of $L^{2}(\mathbb{R}^{2})$ where the impact of the non-symmorphic nature of this $\Gamma$ is illustrated and the twist coming from the vector system is clear.
Acknowledgements: Our interest in finding a characterization of $\Gamma$–shift-invariant closed subspaces originated with a question posed by Azita Mayeli. We also thank a referee of an earlier version of this paper who brought [3] to our attention.
@@ 3 Crystal groups
Fix a cross-section $\gamma:\Pi\to\Gamma$ of the $T$-cosets in $\Gamma$; so $q\big(\gamma(M)\big)=M$, for all $M\in\Pi$. With $\gamma$ fixed, for each $M\in\Pi$, let $x_{M}\in\mathbb{R}^{n}$ be such that $\gamma(M)=[x_{M},M]$. Then $\Gamma=\{[\ell+x_{M},M]:\ell\in\mathcal{L},M\in\Pi\}$. For many crystal groups, $0,M]\in\Gamma$, for all $M\in\Pi$. Then, choose $x_{M}=0$, for each $M\in\Pi$. When this can be done, $\Gamma$ is isomorphic to the semidirect product $\mathcal{L}\rtimes\Pi$. As pointed out in the introduction, such crystal groups are called symmorphic.
Definition 3.1.
Let $\Gamma$ be a crystal group in dimension $n$ with point group $\Pi$. A vector system associated with $\Gamma$ is a map $M\mapsto x_{M}$ of $\Pi$ into $\mathbb{R}^{n}$ such that $x_{M},M]\in\Gamma$, for all $M\in\Pi$.
@@
If $\{x_{M}:M\in\Pi\}$ is a vector system associated with $\Gamma$, then $N^{-1}x_{K}+x_{N}-x_{KN}\in\mathcal{L}$, for all $K,N\in\Pi$.
@@
This auxiliary group might be isomorphic to $\Gamma$, but this is not always the case. As with $\Gamma$, $\mathcal{L}^{*}\rtimes\Pi$ acts on $\mathbb{R}^{n}$. For $(\nu,M)\in\mathcal{L}^{*}\rtimes\Pi$ and $\omega\in\mathbb{R}^{n}$, let $(\nu,M)\cdot\omega=M(\omega+\nu)$. This identifies $\mathcal{L}^{*}\rtimes\Pi$ with a discrete group of isometries of $\mathbb{R}^{n}$ such that, since $\mathcal{L}^{*}$ is a full-rank lattice, $\mathbb{R}^{n}/(\mathcal{L}^{*}\rtimes\Pi)$ is compact. That is, $\mathcal{L}^{*}\rtimes\Pi$ is also a crystal group. Let $\Gamma^{*}=\mathcal{L}^{*}\rtimes\Pi$.
@@ 4 Several Unitary Transformations
Throughout this section, $\Gamma$ is a fixed crystal group with $\mathcal{L}$, $\mathcal{L}^{*}$, $\Pi$, $T$, $\Omega$, and a vector system $\{x_{M}:M\in\Pi\}$ as in Section 3. In order to illuminate aspects of the natural unitary representation of the crystal group $\Gamma$ in the next section, the Hilbert space $L^{2}(\mathbb{R}^{n})$ will be transformed into a number of different realizations by specific unitary maps. The first is the Fourier transform viewed as a unitary map of $L^{2}(\mathbb{R}^{n})$ onto itself.
@@
The function $w_{2}$ plays an important role in the characterization of $\Gamma$–shift-invariant closed subspaces of $L^{2}(\mathbb{R}^{n})$ obtained in Section 6, so it is important to note that this function is independent of the choice of the vector system associated with $\Gamma$. Indeed, if $\{x^{\prime}_{L}:L\in\Pi\}$ is another vector system associated with $\Gamma$, then, for each $L\in\Pi$, $x_{L},L]^{-1}[x^{\prime}_{L},L]=[x^{\prime}_{L}-x_{L},{\rm id}]\in T$, so $k=x^{\prime}_{L}-x_{L}\in\mathcal{L}$. Thus, $e^{-2\pi i\nu\cdot x^{\prime}_{L}}=e^{-2\pi i\nu\cdot(x_{L}+k)}=e^{-2\pi i\nu\cdot x_{L}}$, for all $\nu\in\mathcal{L}^{*}$.
@@ 5 The natural representation
We continue with the notation of the previous section. The action of $\Gamma$ on $\mathbb{R}^{n}$ generates a unitary representation. For $x,M]\in\Gamma$, $\pi[x,M]$ is the unitary operator given by, for $f\in L^{2}(\mathbb{R}^{n})$,
@@ 6 Invariant closed subspaces
The patch of brick wall illustrated in Figure 1 is meant to be a region of a pattern extending in all directions. The symmetry group of this pattern is often denoted $pg$, so set $\Gamma=pg$. If the origin is placed at the bottom of one of the vertical line segments, let $\bf u$} denote a horizontal vector pointing right whose length equals the length of one brick. Let $\bf v$} be an upward pointing vector whose length equals twice the width of a brick. Then, we can take the translation lattice of $\Gamma$ to be $\mathcal{L}=\{j{\bf u}+k{\bf v}:j,k\in\mathbb{Z}\}$. Let $\sigma\in{\rm O}_{2}$ denote the reflection about the horizontal axis. This is not a symmetry of the pattern, but $\frac{1}{2}{\bf u},\sigma]\in\Gamma$. Indeed, the point group of $\Gamma$ is $\Pi=\{{\rm id},\sigma\}$, isomorphic to $\mathbb{Z}_{2}$. If $T=\{[j{\bf u}+k{\bf v},{\rm id}]:j,k\in\mathbb{Z}\}$ is the translation subgroup of $\Gamma$, then $\{{\rm id}\mapsto 0;\ \sigma\mapsto\frac{1}{2}{\bf u}\}$ is a vector system associated with $\Gamma$.
@@
Let’s construct a simple $\Pi$-invariant measurable range function for $\ell^{2}(\mathcal{L}^{*})$ based on $\Pi\Omega$. Let $\kappa=\begin{pmatrix}1\\ 0\end{pmatrix}\in\mathcal{L}^{*}$ and consider the one-dimensional subspace $V=\mathbb{C}(\delta_{0}+\delta_{\kappa})=\{\alpha(\delta_{0}+\delta_{\kappa}):\alpha\in\mathbb{C}\}$ of $\ell^{2}(\mathcal{L}^{*})$. Let $E$ be a measurable subset of $\Omega$ as pictured in Figure 3 where we have suppressed the vertical axis and changed the scale a little.
@@
which by Lemma 6.6 is equivalent to $G(L\omega)\in J^{\prime}(L\omega)$ for a.e. $\omega$, where $J^{\prime}$ is the $\Pi$-invariant extension of $J$. Thus $G\in\mathcal{M}_{J^{\prime}}$. It now follows from Theorem 6.3 that $\mathcal{V}$ is invariant under shifts by $\Gamma$. ∎